OpenStudy (anonymous):

Does -5/9+2/3= -11/9???

OpenStudy (anonymous):

nope it is \[(-5+6)/9=1/9\]

OpenStudy (anonymous):

-(5)/(9)+(2)/(3) To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is 9. Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. (2)/(3)*(3)/(3)-(5)/(9) Complete the multiplication to produce a denominator of 9 in each expression. (6)/(9)-(5)/(9) Combine the numerators of all fractions that have common denominators. (6-5)/(9) Subtract 5 from 6 to get 1. (1)/(9)

OpenStudy (anonymous):

\[\frac{-5}{9}+\frac{2}{3}=\frac{-5}{9}+\frac{6}{9}=\frac{-5+6}{9}=\frac{1}{9}\]

OpenStudy (anonymous):

right kmousou

OpenStudy (anonymous):

great guys, thank you so so much

OpenStudy (anonymous):

our pleasure

OpenStudy (anonymous):

our pleasure